What comes next in the series 15, 13, 10, 8, 5, 3
Answer by stanbon(75887)

What comes next in the series
15
15-2 = 13
13-3 = 10
10-2 = 8
8-3 = 5
5-2 = 3
3-3 = 0
==============
Cheers,
Stan H.
================